# From Canopy to Ground via ForestGen3D: Learning Cross-Domain Generation of 3D Forest Structure from Aerial-to-Terrestrial LiDAR

### v1

@@ 3.6 Discussion
Our findings confirm that DDPM models can generate realistic 3D representations of forest tree structures. The model’s ability to preserve fidelity to the true 3D structure depends on the informativeness of additional conditioning data sources. Furthermore, the computational complexity of the generative process scales linearly with area coverage, enabling efficient deployment across large landscapes without increasing model complexity. With current implementations taking $\approx 2$ seconds per tree-scale generation in a Tesla V100S-PCIE-32GB GPU and with possibilities to further parallelize generation. However, the approach has several limitations. First, the generative model is limited by the distribution it has learned from the training data, which in this case primarily consists of mixed coniferous forest types. When applied to generate trees with structural characteristics not well represented in the training set such as species with drastically different crown architectures, stem densities, branching patterns or ground vegetation (e.g., small shrubs, debri), the model may produce structures that deviate from true vegetation distributions. As a result, generalization to out-of-distribution ecosystems, such as boreal forests, tropical canopies, or hardwood-dominated stands, remains open. Addressing this limitation will require either broadening the training dataset to cover more ecological diversity or incorporating domain adaptation techniques to enable transfer across forest types with different structural priors. While the ALS convex hull provides a useful geometric prior for bounding plausible TLS generation, it is important to note that this containment assumption is approximate. The convex envelope may under/overestimate volume in sparsely sampled regions, and containment does not guarantee structural realism on its own. However, we empirically observe that TLS scans lie within the ALS envelope with over 96% consistency in the training set, and that generation fidelity to the true structure and containment are strongly correlated. These findings support its use as a practical diagnostic metric in large-scale, ground-truth limited deployments.
@@
In dense forest stands, tree detection is often challenged by dense canopies that occlude crown boundaries in ALS data. As a result, the detection pipeline may assign a single bounding box to multiple adjacent trees, which in turn affects the input to the generative model. Depending on the structural cues available in the ALS input, such as crown asymmetry, partial separation, or variations in canopy height the generative model may synthesize either a single fused structure or multiple distinct stems. Because the training dataset includes several examples of closely packed multi-tree configurations, the capacity to distinguish between adjacent trees when sufficient information is present in the point cloud. Nevertheless, in cases where ALS input lacks clear separation cues, the resulting generations may overestimate stem diameter or merge tree forms. To improve robustness in these scenarios, future work will explore both during training and deployment, forest context–aware adaptive tree detection methods, including the use of segmentation masks for individual crown delineation as in [26], as well as stem count aware conditioning strategies to better guide the generative process when multiple trees are present within a single detection. Finally, the model does not explicitly account for temporal variations in forest structure caused by seasonal changes, prescribed burns, wildfires, thinning, or disease. Ignoring these dynamic factors can create a mismatch between generated and real-world forest conditions, limiting the model’s applicability for long-term ecological monitoring and fuel characterization. The aim in future studies is to develop models with additional training/conditional data informative of the potential temporal dynamics either occurring naturally or from human intervention.
Despite these limitations, the ability to reconstruct fine-scale 3D vegetation structure from ALS at the tree, plot and regional scales in mixed conifer ecosystems nonetheless offers immediate utility for existing wildfire modeling frameworks. For example, FastFuels [27] aims to operationalize spatially explicit fuel characterizations by integrating LiDAR and field data through inference-driven mapping. ForestGen3D can complement this by filling in sub-canopy structural gaps where TLS is unavailable, thereby improving vertical fuel profile estimates. Similarly, FIRETEC [28], a physics-based coupled fire–atmosphere model, requires high-resolution 3D inputs that capture canopy gaps, ladder fuels, and understory variation to accurately simulate fire behavior. The generative fidelity of ForestGen3D provides a scalable means to supply such structure across broad landscapes. In addition, QUIC-Fire [29], which enables fast, reduced-order simulations in heterogeneous vegetation, can benefit from ALS-guided 3D generation to populate large domains with spatially consistent and ecologically realistic fuel representations. Integrating ForestGen3D into these modeling frameworks could ultimately support more accurate fire behavior predictions, treatment evaluations, and scenario based risk assessments.
@@ 4 Conclusion
We present ForestGen3D, a scalable generative modeling framework that produces high-resolution 3D forest structure from aerial LiDAR alone. By leveraging denoising diffusion probabilistic models trained on co-registered aerial and terrestrial LiDAR data, ForestGen3D reconstructs sub-canopy detail that is typically occluded in ALS, enabling realistic and ecologically grounded 3D generations across landscapes. Our work introduces a geometric containment prior, referred to as the expected point consistency, preserved over model generations ensuring spatial plausibility, and we show that this prior can serve as a useful proxy for reconstruction quality in the absence of ground truth TLS. Empirical results across multiple spatial scales confirm that ForestGen3D approximates TLS structure with high fidelity in both geometric and ecological terms. Biophysical metrics derived from generated data such as tree height, DBH, crown diameter, and volume closely match those from full ALS+TLS data, demonstrating the model’s utility in ecological monitoring and risk analysis. Despite these strengths, the model’s generalizability is constrained by the structural diversity present in the training dataset. Ecosystems with markedly different vegetation forms or disturbance regimes may require adaptation via domain transfer or broader training sets. In addition, the model currently does not explicitly account for temporal dynamics such as seasonal change, management treatments, or natural disturbance recovery. Future work will extend the framework to incorporate more ecosystem diversity, temporal dynamics, and time-aware conditioning to further improve fidelity and robustness. Overall, ForestGen3D provides a practical and extensible solution to the challenge of generating detailed 3D forest structure in ALS-only environments, with promising applications in ecological forecasting, remote sensing, and wildfire modeling.
@@ Acknowledgement
Research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220024DR. This research was also funded by the Department of Defense, Environmental Security Technology Certification Program, grant number RC24-8162. The findings and conclusions in this publication are those of the authors and should not be construed to represent any official USDA or U.S. Government determination or policy.
Appendix A: Soft Containment Guarantee via ELBO Minimization
Assumptions. Let $(\mathbf{X}_{0},\mathbf{Y})$ denote a paired TLS–ALS point cloud sample drawn from the empirical training distribution $\mathcal{D}$. Let $\mathcal{B}_{\text{ALS}}(\mathbf{Y})=\mathrm{Conv}(\mathbf{Y})$ denote the convex hull of the ALS point cloud. Define the point-level containment indicator $C(\mathbf{x},\mathbf{Y})=1(\mathbf{x}\in\mathcal{B}_{\text{ALS}}(\mathbf{Y}))$.
@@
Figure 14: DDPM model generation under unconditional setting: Sample from Gaussian distribution and generate tree structures according to a training TLS scan dataset.
@@
Appendix C: Additional Regional-scale Generations
This section just presents a few more illustrations of the generation performance of our ForestGen3D model at the regional scale. The included Figures below are spatial snapshots of the 200m radius regions. In these figures, the sources are color-coded as: magenta for ALS, turquoise for TLS, peach for ALS + ForestGen3D generation, and white for the generated ForestGen3D sample alone.
@@
Figure 16: Additional examples of landscape generation with fidelity to ALS scan inputs. Here the ForestGen3D model was targeted only to regions where there is occlusion from vegetation and from an aerial perspective.

### v2

@@ 3.7 Discussion
Our findings confirm that DDPM models can generate realistic 3D representations of forest tree structures. The model’s ability to preserve fidelity to the true 3D structure depends on the informativeness of additional conditioning data sources. Furthermore, the computational complexity of the generative process proposed here scales linearly with area coverage, enabling efficient deployment across large landscapes without increasing model complexity. With current implementations taking $\approx 0.4$ seconds per tree-scale generation in a Tesla V100S-PCIE-32GB GPU and with possibilities to further parallelize generation. However, the approach has several limitations. First, the generative model is limited by the distribution it has learned from the training data, which in this case primarily consists of mixed coniferous forest types. When applied to generate trees with structural characteristics not well represented in the training set such as species with drastically different crown architectures, stem densities, branching patterns or ground vegetation (e.g., small shrubs, debri), the model may produce structures that deviate from true vegetation distributions. As a result, generalization to out-of-distribution ecosystems, such as boreal forests, tropical canopies, or hardwood-dominated stands, remains open. Addressing this limitation will require either broadening the training dataset to cover more ecological diversity or incorporating domain adaptation techniques to enable transfer across forest types with different structural priors. While the ALS convex hull provides a useful geometric prior for bounding plausible TLS generation, it is important to note that this containment assumption is approximate. The convex envelope may under/overestimate volume in sparsely sampled regions, and containment does not guarantee structural realism on its own. However, we empirically observe that TLS scans lie within the ALS envelope with over 96% consistency in the training set, and that generation fidelity to the true structure and containment are strongly correlated. These findings support its use as a practical diagnostic metric in large-scale, ground-truth limited deployments. In addition, while single-scan TLS provides a practical and widely used means of capturing tree-level structure, we recognize that multi-scan, multi-view TLS acquisitions would offer a more complete 3D characterization of vegetation by substantially reducing occlusion, especially in lower and mid-story foliage. Access to such high-quality, fully co-registered TLS datasets would benefit generative modeling in two ways. First, it would provide a less biased training distribution, allowing the DDPM to learn under-story geometry from point clouds that better approximate the true forest structure. Second, improved completeness in the TLS training dataset would likely enhance the model’s ability to reproduce fine-scale geometric features and further improve downstream LiDAR-based biometric estimate tasks. Exploring multi-scan TLS datasets therefore represents a promising direction for future work, particularly for advancing the model’s capacity to learn and generate detailed understory and ground-fuel structure.
@@
In dense forest stands, tree detection is often challenged by dense canopies that occlude crown boundaries in ALS data. As a result, the detection pipeline may assign a single bounding box to multiple adjacent trees, which in turn affects the input to the generative model. Depending on the structural cues available in the ALS input, such as crown asymmetry, partial separation, or variations in canopy height the generative model may synthesize either a single fused structure or multiple distinct stems. Because the training dataset includes several examples of closely packed multi-tree configurations, the capacity to distinguish between adjacent trees when sufficient information is present in the point cloud. Nevertheless, in cases where ALS input lacks clear separation cues, the resulting generations may overestimate stem diameter or merge tree forms. To improve robustness in these scenarios, future work will explore both during training and deployment, forest context–aware adaptive tree detection methods, including the use of segmentation masks for individual crown delineation as in [38], as well as stem count aware conditioning strategies to better guide the generative process when multiple trees are present within a single detection. Finally, the model does not explicitly account for temporal variations in forest structure caused by seasonal changes, prescribed burns, wildfires, thinning, or disease. Ignoring these dynamic factors can create a mismatch between generated and real world forest conditions, limiting the model’s applicability for long-term ecological monitoring and fuel characterization. The aim in future studies is to develop models with additional training/conditional data informative of the potential temporal dynamics either occurring naturally or from human intervention.
Despite these limitations, the ability to reconstruct fine-scale 3D vegetation structure from ALS at tree, plot, and regional scales offers immediate value for next-generation wildfire and ecosystem modeling frameworks. ForestGen3D generates spatially explicit 3D point clouds that preserve the heterogeneous landscape geometry encoded in ALS while reconstructing missing sub-canopy and stem level structure. This capability is particularly important for advancing next-generation fire behavior and fuel modeling systems based on FIRETEC [39], QUIC-Fire [40], and FASTFuels [4], which are beginning to use 3D representations of canopy gaps, ladder fuels, and vertical fuel continuity. By producing high-resolution 3D forest structure that remains spatially consistent with ALS observations, ForestGen3D enables high-resolution 3D models to be applied in ALS-only environments while supplying the sub-canopy detail required for physically realistic simulation. While ALS-based allometric models have proven effective for estimating scalar forest attributes such as DBH, biomass, and crown dimensions from canopy-level measurements, they fundamentally operate on summary metrics and do not reconstruct explicit three-dimensional geometry. In this work, biometric estimation is therefore treated as a downstream evaluation task rather than the primary objective. Our results show that ForestGen3D not only reconstructs realistic 3D structure but also yields biophysical metric distributions that closely match those derived from TLS. Quantitative evaluations using distributional similarity metrics and RMSE demonstrate that ALS+ForestGen3D improves DBH estimation relative to conventional ALS-based allometry, indicating that the model captures ecologically meaningful sub-canopy and stem-level variability.
@@ 4 Conclusion
We present ForestGen3D, a scalable generative modeling framework that produces high-resolution 3D forest structure from aerial LiDAR alone. By leveraging denoising diffusion probabilistic models trained on co-registered aerial and terrestrial LiDAR data, ForestGen3D reconstructs sub-canopy detail that is typically occluded in ALS, enabling realistic and ecologically grounded 3D generations across landscapes. Our work introduces a geometric containment prior, referred to as the expected point consistency, preserved over model generations ensuring spatial plausibility, and we show that this prior can serve as a useful proxy for reconstruction quality in the absence of ground truth TLS. Empirical results across multiple spatial scales confirm that ForestGen3D approximates TLS structure with high fidelity in both geometric and ecological terms. Biophysical metric distributions derived from generated data such as tree height, DBH, crown diameter, and volume closely match those from full ALS+TLS data, demonstrating the model’s utility in ecological monitoring and risk analysis. Despite these strengths, the model’s generalizability is constrained by the structural diversity present in the training dataset. Ecosystems with markedly different vegetation forms or disturbance regimes may require adaptation via domain transfer or broader training sets. In addition, the model currently does not explicitly account for temporal dynamics such as seasonal change, management treatments, or natural disturbance recovery. Future work will extend the framework to incorporate more ecosystem diversity, temporal dynamics, and time-aware conditioning to further improve fidelity and robustness. Overall, ForestGen3D provides a practical and extensible solution to the challenge of generating detailed 3D forest structure in ALS only environments, with promising applications in ecological forecasting, remote sensing, and wildfire modeling.
@@ Acknowledgement
Research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220024DR and 20260010DR. This research was also funded by the Department of Defense, Environmental Security Technology Certification Program, grant number RC24-8162. The findings and conclusions in this publication are those of the authors and should not be construed to represent any official USDA or U.S. Government determination or policy.
Appendix A: Expected Point Consistency Guarantee via ELBO Minimization
Assumptions. Let $(\mathbf{X}_{0},\mathbf{Y})$ denote a paired TLS/ALS point cloud sample drawn from the empirical training distribution $\mathcal{D}$. Let $\mathcal{B}_{\text{ALS}}(\mathbf{Y})=\mathrm{Conv}(\mathbf{Y})$ denote the convex hull of the ALS point cloud. Define the point-level containment indicator $C(\mathbf{x},\mathbf{Y})=1(\mathbf{x}\in\mathcal{B}_{\text{ALS}}(\mathbf{Y}))$.
@@
Figure 14: ForestGen3D
model generation under the unconditional setting: Sample from Gaussian distribution and generate tree structures according to a training TLS scan dataset.
@@
Appendix C: Additional comparisons of biophysical metric distributions
Figures 15, 16, and 17 present additional plot-scale examples (25 m radius) of biophysical metric distributions for tree height, DBH, crown diameter, and crown volume, derived from ALS, TLS, ALS+TLS, and ALS+ForestGen3D point clouds. Across all examples, ALS+ForestGen3D consistently compensates for ALS structural blind spots while preserving ALS-derived spatial geometry. Tree height and crown diameter distributions remain largely contained within ALS-derived ranges, whereas the most pronounced improvements are observed for DBH and crown volume, reflecting recovery of sub-canopy and lower-canopy structure absent in ALS-only data. Collectively, these results further confirm that ForestGen3D produces structurally plausible 3D vegetation whose biometric distributions closely align with TLS-derived references while maintaining fidelity to ALS-provided canopy geometry.
Figure 15: Plot-scale distributions of biophysical tree metrics derived from different sensing sources. Rows correspond to ALS (row 1), TLS (row 2), ALS+TLS (row 3), and ALS+ForestGen3D (row 4). Columns show distributions for tree height (col. 1), diameter at breast height (DBH; col. 2), crown diameter (CrD; col. 3), and crown volume (CrV; col. 4). Each histogram represents metrics computed over a 25 m radius plot. ALS+TLS is treated as the structural reference, while ALS+ForestGen3D closely approximates its metric distributions, indicating successful recovery of sub-canopy structure while preserving ALS-derived spatial constraints.
Figure 16: Plot-scale distributions of biophysical tree metrics derived from different sensing sources. Rows correspond to ALS (row 1), TLS (row 2), ALS+TLS (row 3), and ALS+ForestGen3D (row 4). Columns show distributions for tree height (col. 1), diameter at breast height (DBH; col. 2), crown diameter (CrD; col. 3), and crown volume (CrV; col. 4). Each histogram represents metrics computed over a 25 m radius plot. ALS+TLS is treated as the structural reference, while ALS+ForestGen3D closely approximates its metric distributions, indicating successful recovery of sub-canopy structure while preserving ALS-derived spatial constraints.
Figure 17: Plot-scale distributions of biophysical tree metrics derived from different sensing sources. Rows correspond to ALS (row 1), TLS (row 2), ALS+TLS (row 3), and ALS+ForestGen3D (row 4). Columns show distributions for tree height (col. 1), diameter at breast height (DBH; col. 2), crown diameter (CrD; col. 3), and crown volume (CrV; col. 4). Each histogram represents metrics computed over a 25 m radius plot. ALS+TLS is treated as the structural reference, while ALS+ForestGen3D closely approximates its metric distributions, indicating successful recovery of sub-canopy structure while preserving ALS-derived spatial constraints.
Appendix D: Additional Regional-scale Generations
This section provides additional visual examples of ForestGen3D outputs at the regional scale. Figures 18 and 19 show spatial snapshots of 200 m radius regions, with point clouds color-coded as follows: magenta for ALS, turquoise for TLS, peach for ALS combined with ForestGen3D output, and white for ForestGen3D-generated points alone. Note that ForestGen3D is applied only within regions identified as occluded in the ALS data. Thus, figures displaying only the generated points contain empty areas corresponding to locations where no synthesis was performed. When combined with the original ALS observations, these regions are filled by existing ALS ground and canopy returns, resulting in a complete point cloud representation. The generated points primarily populate the lower canopy and sub-canopy layers and extend toward the ground surface, following the spatial constraints imposed by the ALS geometry. The synthesized structures exhibit continuous transitions between ALS-observed and generated regions, yielding coherent tree-level and stand-level configurations. Each 200 m radius region contains approximately 3,000 trees. With an average inference time of about 0.4 seconds per tree, generation of a full region requires on the order of 20 minutes, illustrating the computational feasibility of applying ForestGen3D at regional scales.
@@
Figure 18: Additional examples of landscape-scale ForestGen3D generation. Each row corresponds to a distinct 200 m radius area centered on a TLS scan location. Columns show (left) ALS combined with TLS reference data, (middle) ALS combined with ForestGen3D output, and (right) ForestGen3D-generated points alone. Point clouds are color-coded as follows: magenta for ALS, turquoise for TLS, peach for ALS + ForestGen3D, and white for ForestGen3D-generated points. ForestGen3D is applied selectively to regions exhibiting occlusion in the ALS data, generating sub-canopy structure while preserving ALS-observed geometry.
@@
Figure 19: Additional landscape-scale ForestGen3D examples. Rows show different 200 m radius areas; columns show ALS+TLS (left), ALS+ForestGen3D (middle), and ForestGen3D alone (right). Colors denote ALS (magenta), TLS (turquoise), ALS+ForestGen3D (peach), and generated points (white).